A printing order for 49920 good flyers is received. If 2.5% of the flyers that are printed will be discarded because of defects, how many flyers should be printed to assure you have 49920 good ones?
Answer by TimothyLamb(4379)

p * (1 - 0.025) = 49920
p * (0.975) = 49920
p = 49920/0.975
p = 51200
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php